State whether the following statements are true or false Give reasons for your answers.
1. The function f:R^3->R, is given by f(x,y,z)=|x|+|y|+|z| is differentiable at (2,3,-1)
2. The function f(x,y)= max{y/x,X} is a homogeneous function on R^2.
3. The domain of the function f / g where f(x,y)=2xy and g(x,y)=x^2+y^2 is R^2.
1
Expert's answer
2020-05-15T17:47:24-0400
  1. True. In small neighbourhood of "p=(2,3,-1)" we have "f(x,y,z)=x+y-z" that is obviously differentiable in the neighbourhood of "p".
  2. False. Let "a=-1", then "\max\{ay/x, ax\}\ne a\max\{y/x, x\}". For example, "x=1,y=7\colon \max \{-7,-1\}\ne-\max\{7,1\}=-7"
  3. False. Domain is "\{(x,y)\in\mathbb R^2\colon x^2+y^2>0\}". Particularly, "\lim\limits_{(x,y)\to(0,0)} \frac{f(x)}{g(x)}" does not exist because for "(x,y)=(1\/n,k\/n)" we have different limits "\lim\limits_{n\to \infty}\frac{f}{g}=\frac{2k}{1+k^2}" for different "k"